### v2

# DPLib: A Standard Benchmark Library for Distributed Power System Analysis and Optimization

Abstract
DPLib is an open-source MATLAB-based benchmark library created to support research and development in distributed and decentralized power system analysis and optimization. Distributed and decentralized methods offer scalability, privacy preservation, and resilience to single points of failure, making them increasingly important for modern power systems. However, unlike centralized tools such as MATPOWER, no general-purpose, reproducible data library package currently exists for distributed power system studies. DPLib, available at GitHub, fills this gap by providing a standard power system library featuring over 20 multi-region benchmark test cases of varying sizes, along with a graph-based partitioning toolkit that decomposes any MATPOWER test system into multiple electrically coherent regions. The partitioning toolkit, an easy-to-use MATLAB code, generates standardized .mat and .m files, along with region visualizations for intuitive understanding. We also provide modular, easy-to-use distributed optimal power flow (OPF) solvers: an alternating direction method of multipliers(ADMM)-based DC-OPF solver implemented in YALMIP, and an ADMM-based AC-OPF solver leveraging IPOPT. These solvers validate the generated test systems for distributed optimization applications. Numerical results validate the generated test cases, establishing DPLib as a foundation for reproducible distributed power system research.
Index Terms:
Distributed power grid test systems, standard benchmark library, MATPOWER, distributed optimization.
I Introduction
Centralized frameworks have long served as the dominant approach for modeling and solving large-scale power system problems, including optimal power flow (OPF) and security-constrained unit commitment. A key enabler behind the success and widespread adoption of centralized optimization and analysis in power systems is the availability of standardized benchmark test cases and libraries. Among these, MATPOWER has emerged as a widely accepted benchmark tool, offering consistent, well-structured data for centralized OPF studies  [1]. Its role in enabling reproducibility, fair performance comparisons, and methodological development has been fundamental to the evolution of centralized power system optimization and analysis.
For centralized OPF studies, for instance, [2] proposed convex relaxations for AC OPF and validated their results on MATPOWER test cases. [3] provided a comprehensive review of relaxation techniques for OPF, with all benchmarks sourced from MATPOWER. [4] introduced QC relaxation for OPF using MATPOWER data. More recently, [5] developed robust optimization methods for corrective control under uncertainty, relying on MATPOWER cases for validation. [6] proposed the NESTA archive to enrich MATPOWER with more realistic data, yet retained full MATPOWER compatibility. These works and many more underscore the importance of a common, reliable data standard—facilitated by MATPOWER—that allows researchers to validate and compare centralized optimization algorithms under consistent and reproducible conditions.
In the last decade, distributed and decentralized computation has gained significant momentum in the power systems community, especially in applications such as distributed OPF, voltage control, state estimation, and demand-side management [7, 8, 9]. These include studies that apply alternating direction method of multipliers (ADMM) decomposition [8, 10], distributed voltage control [11], parallel and distributed OPF [12, 13], consensus-based optimization [14], and cooperative control frameworks [15, 16]. Additional contributions include scalable distributed frameworks for transmission networks [17], graph-based coordination strategies [7], and distributed control in microgrids [18]. These methods offer inherent advantages in scalability, privacy preservation, and resilience to single points of failure.
However, as the literature on distributed techniques expands, a critical gap remains: the lack of unified, publicly available benchmark test cases tailored for distributed power system studies. Without such a standard library, researchers must often construct ad hoc datasets or apply arbitrary partitioning schemes, which hampers reproducibility and limits the comparability of different distributed algorithms. There is a strong need for a publicly available and community-accepted library of distributed test cases that allows researchers to validate, compare, and benchmark their distributed optimization or control methods under a common, well-defined framework.
In response to this need, we develop a distributed power system library, DPLib, that addresses the lack of standardized resources for distributed power system studies by providing a comprehensive library of over 20 benchmark test cases, ranging from 73 to 9241 buses. These multi-region systems are generated using the DPLib graph-based partitioning toolkit that ensures electrical coherence and captures the corresponding multi-regional network topology. Each case includes the necessary structural and connectivity data to support distributed analysis. All test cases have been verified using distributed OPF problems to demonstrate their practicality and confirm their suitability for reproducible and scalable distributed optimization.
In addition to a standard library, a key requirement is a general and user-friendly partitioning tool capable of processing any standard MATPOWER case and dividing it into electrically coherent subregions based on user-defined parameters, such as the number of regions. An effective tool should preserve the integrity of the original network, identify and document inter-regional tie-lines, and produce output in a structured format consistent with centralized MATPOWER cases. This functionality enables users to generate customized test cases while maintaining compatibility with a standard data structure. The DPLib graph-based partitioning toolkit fulfills all of these requirements.
While the distributed library includes several pre-partitioned and validated datasets, researchers may also wish to generate and test their customized partitions using the DPLib graph-based toolkit. To support this, DPLib provides ADMM-based solvers for both DC OPF and AC OPF that are fully compatible with the partitioned data structure produced by the tool. These solvers take the partitioned datasets as input, iteratively solve the regional subproblems, enforce tie-line consistency through Lagrangian multipliers, and monitor convergence metrics, enabling reliable validation of user-defined partitions.
The main contribution of this work is the introduction of DPLib, a unified framework that establishes the first standard distributed data library for power system applications. Besides the library, DPLib consists of three core tools: a graph-based partitioning tool that decomposes any MATPOWER case into multiple regions using spectral clustering and k-means, enabling the creation of distributed datasets; a modular ADMM-based distributed DC OPF solver implemented in YALMIP; and a distributed AC OPF solver using ADMM and IPOPT. The DC OPF and AC OPF solvers are designed to validate and benchmark the partitioned datasets under realistic operating conditions. Together, these tools support reproducibility, scalability, and algorithm testing for distributed power system applications, addressing a key gap in current research infrastructure.
The remainder of the paper is organized as follows. Section II introduces the graph-based modeling of power networks and details the DPLib partitioning toolkit. Section III describes the ADMM-based DC and AC OPF solvers used to validate partitioned datasets. Section IV presents the benchmark case studies generated and tested using DPLib. Finally, Section V concludes the paper.
II Graph-Based Partitioning and Toolkit Description
In many power systems studies, realistic standard test cases in the MATPOWER format are commonly used for testing and validation. This practice ensures that developed methods are evaluated on realistic cases that reflect the complexity of actual power systems. To maintain this standard, researchers working on distributed algorithms should also rely on realistic standard test systems and datasets rather than arbitrarily constructed or manually partitioned grids. In practice, real power networks often exhibit a naturally regional structure. For example, a network covering several regions will typically have strong internal connectivity within each region and only a few tie-lines connecting areas to one another. This results in regions that are strongly connected internally but weakly connected to the other areas. A proper partitioning method should be able to observe these loosely connected boundaries. We use graph-based partitioning to provide a principled and effective framework for capturing such structures and dividing the network in a way that aligns with real systems’ physical and topological characteristics.
II-A Power Network Modeling and Spectral Clustering
To enable distributed optimization and control, a centralized network must first be divided into smaller regions that remain physically meaningful and electrically coherent. This partitioning process starts by modeling the power system as an undirected graph. We represent the power network as $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where:
  • $\mathcal{V}$ is the set of buses (nodes), as defined in the original MATPOWER case file.
  • $\mathcal{E}$ is the set of transmission lines (edges), each connecting two buses.
The graph is undirected because the power flow direction does not impact the structure for partitioning purposes—only a connection between buses matters.
A path in a graph is a sequence of nodes such that an edge connects each consecutive pair of nodes in the sequence. A connected graph is an undirected graph with a path between every pair of nodes. This means it is possible to travel from any node to any other by following a sequence of edges in the graph. To describe the structure of $\mathcal{G}$, we construct the adjacency matrix $A\in\mathbb{R}^{n\times n}$, where $n=|\mathcal{V}|$ is the total number of buses. The entries of this matrix are defined as:
$A_{ij}=\begin{cases}1,&\text{if there is a line connecting bus }i\text{ and bus }j,\\ 0,&\text{otherwise}.\end{cases}$
Because the graph is undirected, $A$ is symmetric. Next, we define the degree matrix $D\in\mathbb{R}^{n\times n}$, a diagonal matrix where each diagonal entry $D_{ii}$ equals the number of lines connected to bus $i$:
$D_{ii}=\sum_{j=1}^{n}A_{ij}.$
Using these two matrices, we define the graph Laplacian $L=D-A.$ The Laplacian matrix $L$ encodes the power network connectivity and topology. For a connected graph, $L$ is symmetric and positive semi-definite, meaning all its eigenvalues are non-negative. Importantly, the smallest eigenvalue of $L$ is always zero, corresponding to the trivial uniform eigenvector.
To divide the system into $k$ electrically coherent regions, we use spectral clustering, which leverages the properties of $L$. The first step is to compute the eigenvalues and eigenvectors of $L$. Among these, the $k$ smallest nonzero eigenvalues are selected. Each eigenvalue is associated with an eigenvector that varies smoothly across the graph, meaning that buses that are strongly connected in the physical system will have similar eigenvalues. Therefore, these eigenvectors reflect the underlying structure of the network.
We form a matrix $U\in\mathbb{R}^{n\times k}$ by placing the $k$ selected eigenvectors as columns. Each row $u_{i}\in\mathbb{R}^{k}$ corresponds to a bus and represents it in a new $k$-dimensional spectral space, where geometrical proximity reflects topological connectivity.
To assign buses into one of $k$ regions, we apply the k-means clustering algorithm to the rows of $U$. This unsupervised learning algorithm operates by:
  • Randomly initializing $k$ centroids (in our case, from the rows of $U$);
  • Assigning each bus to the nearest centroid based on Euclidean distance;
  • Recalculating the centroid of each cluster;
  • Repeating the assignment and centroid update steps until convergence.
The clustering result yields a partition of all buses into $k$ regions. Each cluster groups strongly connected buses, leading to electrically coherent and balanced regions. After forming the clusters, inter-regional tie-lines (lines that connect buses belonging to different clusters) are identified. To improve quality, the algorithm may repeat the clustering multiple times and select the configuration that minimizes the number of inter-regional tie-lines.
This spectral clustering approach provides a systematic and scalable method for decomposing any MATPOWER case into meaningful regions. It aligns well with practical expectations since most real-world partitioned power networks are formed based on geographic or electrical proximity.
A tool that can automatically convert MATPOWER test cases into $k$ well-structured regions is essential for distributed power system studies. The DPLib partitioning toolkit fills this gap by generating standardized multi-region datasets from any MATPOWER case, supporting reproducibility and scalability.
II-B DPLib Partitioning Toolkit
The DPLib graph-based partitioning toolkit implements the method described above and automates the generation of partitioned MATPOWER cases. The partitioning toolkit is publicly available at GitHub, including the MATLAB implementation file named partitioning_code.m. The code uses MATLAB R2020a or newer, along with the MATPOWER toolbox. Upon running the script, the user is prompted to enter the name of a valid MATPOWER case file (for example, case118) and specify the desired number of regions into which the system should be divided.
Internally, the code loads the MATPOWER case and checks whether the bus numbers are sequential. If not, it renumbers the buses from one to $n$ and updates all corresponding references in the branch and generator matrices. It then constructs the undirected adjacency matrix based on the network’s branch connectivity and forms the corresponding graph Laplacian. The eigenvectors of the Laplacian are computed and used as features for $k$-means clustering. The clustering process is repeated multiple times to ensure optimality, and the configuration with the lowest number of inter-regional tie-lines is selected.
For each candidate partitioning, regional subcases are formed by extracting internal buses, branches, generators, and generation cost data for each region. The script ensures that each region has at least one generator. Buses are then renumbered locally within each region to maintain consistency. Inter-regional tie-lines are identified as any branch connecting buses in different regions and stored.
Once a feasible partitioning is found, the code saves the result in multiple formats. A MATLAB .mat file is created, containing all regional subcases under separate variables (e.g., mpc_regionR1, mpc_regionR2, etc.), along with a comprehensive table of all inter-regional tie-lines labeled by their originating and receiving regions and complete branch data. A corresponding .m file is also generated, exporting the same structures in a human-readable script format suitable for inspection or direct use in other codes. Additionally, the code creates a graphical visualization of the partitioned system, where each region is displayed as a labeled node in a circular layout, and tie-lines are represented as edges connecting the areas. The number of tie-lines between each pair of regions is also annotated.
During execution, the code prints detailed messages to the MATLAB command window. These include whether bus renumbering was necessary, confirmation of full graph connectivity, the number of generated regions and tie-lines, the region assignment of the reference (slack) bus, and whether all buses, branches, and generators are fully included in the partitioned result. It also confirms the successful saving of the generated .mat, .m, and .png files. The algorithm of the partitioning toolkit, illustrating the complete data processing pipeline, is provided in 1.
Algorithm 1 DPLib partitioning toolkit
1:User Input:
  • Name of the original MATPOWER case (e.g., case118)
  • Number of desired regions $k$
2: Load the MATPOWER case file
3: Check bus numbering; if non-sequential, renumber from 1 to $n$ and update all references
4: Construct adjacency matrix $A$ from the branch data.
5: Construct degree matrix $D$ with $D_{ii}=\sum_{j}A_{ij}$
6: Form graph Laplacian $L=D-A$
7: Compute the first $k$ nonzero eigenvectors of $L$
8: Stack eigenvectors into matrix $U\in\mathbb{R}^{n\times k}$
9: Apply $k$-means clustering to the rows of $U$
10: Repeat clustering for multiple iterations to minimize the number of tie-lines
11:for each region (cluster) do
12:  Extract internal buses, branches, generators, and cost data
13:  Ensure that at least one generator exists in each region
14:  Locally renumber buses
15:  Identify inter-regional tie-lines
16:end for
17: Save output in the following formats:
  • .mat file with region-wise MATPOWER data
  • .m script file exporting the same data
  • .png visualization showing regions and tie-lines
III Validating DPLib Test Systems Using OPF
To ensure that the generated multi-region test cases are electrically coherent and suitable for distributed power system studies, it is essential to validate them through power system analysis. Although the DPLib partitioning toolkit generates distributed variants of MATPOWER test systems applicable to a wide range of power system studies, including control, state estimation, or stability analysis, this section evaluates these test systems for distributed OPF applications. OPF has been selected as a representative benchmarking problem because of its maturity, availability of centralized references, and compatibility with distributed solvers due to over a decade of research in this area. We have developed general ADMM-based solvers for DC and AC OPF problems that operate directly on the test cases produced by the DPLib partitioning toolkit. These solvers are designed to validate the feasibility, consistency, and convergence of partitioned datasets, demonstrating the suitability of these test systems. Nevertheless, users may use alternative solvers or utilize these provided tools to verify and explore their own partitioned datasets.
III-A Distributed OPF Formulation
We formulate the OPF problem in both DC and AC settings for each region $r$. Each region solves a local OPF problem while maintaining consistency at the boundaries via duplicated (virtual) variables and tie-line constraints. These virtual variables (buses) enable each region to model tie-line interactions locally without requiring direct enforcement of global constraints. The overall structure is illustrated in Fig. 1(a), where a two-region power grid is shown with inter-regional tie-lines in red. Fig. 1(b) further illustrates how virtual buses (shown in red) are added to represent shared variables.
(a) Grid with two regions and red lines are inter-regional transmission tie-lines
(b) Decomposition into two regions with virtual buses shown in red
Figure 1: Illustration of a partitioned power system
III-A1 DC OPF Formulation
The DC OPF for region $r$ is:
$\min_{\boldsymbol{x}^{r}}\sum_{i\in\Omega^{r}_{g}}f(p^{g}_{i})$ (1a)
s.t.
$\qquad\quad\underline{p}_{i}^{g}\leqslant p^{g}_{i}\leqslant\bar{p}_{i}^{g},\quad\forall i\in\Omega^{r}_{g}$ (1b)
$\qquad\quad p^{g}_{i}-p^{d}_{i}=\sum_{j\in\Omega^{r}_{i}}\frac{\theta_{i}-\theta_{j}}{x_{ij}}+\sum_{j\in\Omega^{r,\textrm{in}}_{i}}\frac{\theta_{i}-\tilde{\theta}_{j}}{x_{ij}},\quad\forall i\in\Omega^{r}_{b}$ (1c)
$\qquad\quad\underline{p}_{ij}\leqslant p_{ij}\leqslant\bar{p}_{ij},\quad\forall ij\in\Omega^{r}_{l}\cup\Omega^{r,\textrm{in}}_{l}$ (1d)
$\qquad\quad\theta_{\text{ref }}=0,$ (1e)
$\qquad\quad\theta_{i}=\tilde{\theta}_{i},\quad\tilde{\theta}_{j}=\theta_{j},\quad\forall ij\in\Omega^{r,\textrm{in}}_{l}$ (1f)
$\boldsymbol{x}^{r}=[\boldsymbol{p}^{g\top},\boldsymbol{\theta}^{\top}]\in\mathbb{R}^{n^{r}_{g}+n^{r}_{b}}$ is the decision vector for region $r$, including generation $p^{g}_{i}$ and voltage angles $\theta_{i}$. Equation (1a) minimizes the local cost $f(p^{g}_{i})$. Constraints (1b)–(1c) enforce generator limits and power balance using demand $p^{d}_{i}$, line reactance $x_{ij}$, and virtual angles $\tilde{\theta}_{j}$. Constraint (1d) bounds internal and tie-line flows $p_{ij}$, and (1f) ensures consistency between real and virtual bus angles.
III-A2 AC OPF Formulation
The AC OPF for region $r$ is:
$\min_{\boldsymbol{x}^{r}}\sum_{i\in\Omega^{r}_{g}}f(p^{g}_{i})$ (2a)
s.t.
$\qquad\underline{p}_{i}^{g}\leqslant p^{g}_{i}\leqslant\bar{p}_{i}^{g},\quad\underline{q}_{i}^{g}\leqslant q^{g}_{i}\leqslant\bar{q}_{i}^{g},\quad\forall i\in\Omega^{r}_{g}$ (2b)
$\qquad p^{g}_{i}-p^{d}_{i}=\sum_{j\in\Omega^{r}_{i}}\left(p_{ij,f}+p_{ij,t}\right)+\sum_{j\in\Omega^{r,\text{in}}_{i}}\tilde{p}_{ij},\quad\forall i\in\Omega^{r}_{b}$ (2c)
$\qquad q^{g}_{i}-q^{d}_{i}=\sum_{j\in\Omega^{r}_{i}}\left(q_{ij,f}+q_{ij,t}\right)+\sum_{j\in\Omega^{r,\text{in}}_{i}}\tilde{q}_{ij},\quad\forall i\in\Omega^{r}_{b}$ (2d)
$\qquad p_{ij,f}=G_{ij,1}v_{i}^{2}+G_{ij,2}v_{i}v_{j}\cos(\theta_{i}-\theta_{j})$
$\qquad\qquad\quad+B_{ij,2}v_{i}v_{j}\sin(\theta_{i}-\theta_{j}),\quad\forall ij\in\Omega^{r}_{l}$ (2e)
$\qquad q_{ij,f}=-B_{ij,1}v_{i}^{2}-B_{ij,2}v_{i}v_{j}\cos(\theta_{i}-\theta_{j})$
$\qquad\qquad\quad+G_{ij,2}v_{i}v_{j}\sin(\theta_{i}-\theta_{j}),\quad\forall ij\in\Omega^{r}_{l}$ (2f)
$\qquad p_{ij,t}=G_{ij,4}v_{j}^{2}+G_{ij,3}v_{i}v_{j}\cos(\theta_{j}-\theta_{i})$
$\qquad\qquad\quad+B_{ij,3}v_{i}v_{j}\sin(\theta_{j}-\theta_{i}),\quad\forall ij\in\Omega^{r}_{l}$ (2g)
$\qquad q_{ij,t}=-B_{ij,4}v_{j}^{2}-B_{ij,3}v_{i}v_{j}\cos(\theta_{j}-\theta_{i})$
$\qquad\qquad\quad+G_{ij,3}v_{i}v_{j}\sin(\theta_{j}-\theta_{i}),\quad\forall ij\in\Omega^{r}_{l}$ (2h)
$\qquad p_{ij,f}^{2}+q_{ij,f}^{2}\leq\bar{p}_{ij}^{2}+\bar{q}_{ij}^{2},\quad\forall ij\in\Omega^{r}_{l}\cup\Omega^{r,\textrm{in}}_{l}$ (2i)
$\qquad p_{ij,t}^{2}+q_{ij,t}^{2}\leq\bar{p}_{ij}^{2}+\bar{q}_{ij}^{2},\quad\forall ij\in\Omega^{r}_{l}\cup\Omega^{r,\textrm{in}}_{l}$ (2j)
$\qquad\underline{\Delta\theta}_{ij}\leq\theta_{i}-\theta_{j}\leq\overline{\Delta\theta}_{ij},\quad\forall ij\in\Omega^{r}_{l}\cup\Omega^{r,\textrm{in}}_{l}$ (2k)
$\qquad\theta_{\text{ref}}=0$ (2l)
$\qquad\theta_{i}=\tilde{\theta}_{i},\quad\tilde{\theta}_{j}=\theta_{j},\quad\forall ij\in\Omega^{r,\textrm{in}}_{l}$ (2m)
$\qquad v_{i}=\tilde{v}_{i},\quad\tilde{v}_{j}=v_{j},\quad\forall ij\in\Omega^{r,\textrm{in}}_{l}$ (2n)
The decision vector is $\boldsymbol{x}^{r}=[\boldsymbol{p}^{g\top},\boldsymbol{q}^{g\top},\boldsymbol{\theta}^{\top},\boldsymbol{v}^{\top}]$, including generation $(p^{g}_{i},q^{g}_{i})$, angles $\theta_{i}$, and voltages $v_{i}$ at region $r$. Constraints (2b)–(2d) enforce generator and nodal power limits; (2e)–(2h) define AC flows using admittance terms $G_{ij,k},B_{ij,k}$. Limits on flows and angle differences are imposed in (2i)–(2k). Reference angle (2l) and virtual bus equalities (2m)–(2n) ensure boundary consistency.
III-A3 ADMM Coordination
To ensure consistency between neighboring regions, we use ADMM to coordinate virtual variables [10]. For example, the voltage angle consistency is enforced via:
$L_{\rho}(\theta,\tilde{\theta},\lambda)=f(p^{g})+\lambda^{\top}(\theta-\tilde{\theta})+\frac{\rho}{2}\|\theta-\tilde{\theta}\|^{2},$ (3)
where $\lambda$ are the dual variables and $\rho$ is the penalty parameter. Similar terms are added for voltage magnitudes in AC OPF.
Each ADMM iteration proceeds as:
  1. 1.
    Solve each region’s local OPF using the current dual variables.
  2. 2.
    Update dual variables based on tie-line consistency.
  3. 3.
    Check primal residuals and stop if below a threshold.
In general, coordination across regions is achieved using the ADMM. The distributed DC OPF and AC OPF formulations in (1) and (2) incorporate local objectives and tie-line consistency via augmented Lagrangian. Subproblems are solved locally at each iteration, dual variables are updated, and convergence is checked via the worst primal residual.
III-B General ADMM-Based OPF Solvers
We have developed two open-source general-purpose solvers for distributed OPF problems, one for DC OPF and another for AC OPF. Both solvers adopt ADMM to coordinate regional solutions and ensure consistency across tie-lines.
III-B1 DPLib DC OPF Solver
The DPLib DC OPF solves the OPF problem using ADMM by optimizing each region independently and enforcing agreement on shared variables (tie-line voltage angles) through iterative dual updates. The solver is available at GitHub, including the MATLAB implementation file named DistDCOPF.m. Designed for partitioned MATPOWER cases, it requires MATLAB R2020a or newer with MATPOWER and YALMIP installed. Upon execution, the user provides the partitioned .mat file, ADMM penalty parameter $\rho$, convergence threshold, maximum number of iterations, and optionally the centralized OPF cost for benchmarking via optimality gap.
After receiving user inputs, the code loads all regional and tie-line data, augments each region’s branch matrix with inter-regional tie-lines, and initializes bus voltage angle variables and indexing. The core ADMM loop then solves each region’s DC OPF using YALMIP, minimizing generation cost under local constraints (e.g., power balance, generator limits) while adding consistency penalties on shared tie-line angles as modeled in (1). A helper function handles the regional problem formulation and solution.
After each iteration, the code evaluates tie-line angle consistency and updates dual variables accordingly. Primal residuals are recorded, and convergence is declared if the maximum error drops below the user-defined threshold; otherwise, the process continues until the iteration limit. The optimality gap is computed and plotted alongside residuals if a centralized cost is provided. Progress and convergence status are printed in the command window.
The DPLib DC OPF solver supports any number of regions and is tailored for use with partitioned .mat files, enabling reproducible distributed OPF testing. Algorithm 2 summarizes its workflow.
Algorithm 2 DPLib OPF solver
1:Input: Partitioned MATPOWER data, ADMM parameters $\rho$, convergence tolerance $\epsilon$, maximum iterations
2: Load the partitioned case data for all regions
3: Augment each region with its corresponding tie-line data
4: Initialize region variables including local decision variables and dual variables
5: Retrieve ADMM parameters and consistency constraints
6: Identify and prepare interface buses at region boundaries
7:for each ADMM iteration do
8:  Solve local OPF for each region independently
9:  Update dual variables based on consistency violations at bordering buses
10:  if convergence criteria are met then
11:   break
12:  end if
13:end for
14: Visualize and save final solution
III-B2 DPLib AC OPF Solver
The DPLib AC OPF solver implements a distributed ADMM-based solver in MATLAB for any power system preprocessed using the DPLib partitioning tool. It enables large-scale distributed analysis, where each region independently solves its local AC OPF while coordinating inter-regional tie-lines. The solver developed in this work is publicly available at GitHub, including the MATLAB implementation file named DistACOPF.m. To run the code, the user must have MATLAB R2020a or later, with MATPOWER and IPOPT properly configured. Upon launch, the code prompts for the partitioned .mat file, ADMM penalty parameter $\rho$, convergence threshold, maximum number of iterations, and optionally the centralized OPF cost for benchmarking via optimality gap.
After loading the partitioned data, the code augments each region’s branch matrix with its tie-lines and initializes region-specific variables (e.g., bus counts, initial voltage values). The main ADMM loop iteratively solves each region’s local AC OPF via the helper function Acopf, based on (2), updating voltage magnitudes and angles. Dual variables are adjusted using consistency gradients to enforce agreement across tie-lines. Convergence is declared if the worst primal residual falls below the threshold. The optimality gap is computed per iteration if a centralized cost is provided. Final plots show convergence trends, and status messages are printed throughout. The overall process flow of the DPLib AC OPF solver is presented in algorithm 2.
IV DPLib: Benchmark Test Cases
Distributed versions of many widely used MATPOWER benchmark systems are generated. Each test case is processed using the DPLib graph-based partitioning tool and tested using the DPLib ADMM-based DC and AC OPF solvers described earlier. Across all systems, both solvers successfully converge to feasible solutions. All partitioned cases and solver outputs are publicly available as part of the DPLib at GitHub. In the following subsections, we explain some of the distributed test cases generated and tested as benchmarks in DPLib.
In addition to evaluating convergence using the worst-case primal residual and the optimality gap, further validation tests are performed on the generated multi-region systems to verify their consistency with physical laws and operational constraints of the power grid. Specifically, we verified that upon the convergence of the ADMM iterations, the total power generation across all regions closely matches the total system demand, accounting for minor discrepancies due to modeling approximations and convergence thresholds. Furthermore, we confirmed that active power flows across each inter-regional tie-line are consistent from both sides of the interface. All tested datasets successfully passed these validation checks, further demonstrating the correctness and reliability of the generated partitioned cases.
To facilitate evaluation of the benchmark test cases using OPF solvers, all primal variables were initialized with cold-start conditions: voltage angles were set to zero and voltage magnitudes to one per unit. The dual variables, corresponding to Lagrange multipliers associated with consistency constraints, were initialized to one. Penalty parameters $\rho$ in the ADMM-based solvers were selected case by case to ensure convergence.
IV-A Case: pglib_opf_case200_tamu
The pglib_opf_case200_activ system is a synthetic 200-bus benchmark from the IEEE PES Power Grid Library for optimal power flow studies. Geographically situated in the central region of Illinois, this case models a 230/115 kV transmission network designed to serve a load distribution that mimics the population profile of its geographic footprint. The transmission system was generated following the methods outlined in [19], ensuring statistical similarity to real-world grids without modeling any actual transmission lines. Developed as part of the U.S. ARPA-E GRID DATA research project, the system contains no critical energy infrastructure information and was created using PowerWorld Simulator, then converted to MATPOWER format using version 6. The dataset is publicly available through PGLib and licensed under the Creative Commons Attribution 4.0 International License.
The pglib_opf_case200_tamu test case was partitioned using our graph-based spectral clustering method, which ensured full bus inclusion and topological connectivity. The best clustering configuration identified by the algorithm consisted of three electrically coherent regions, interconnected through 10 tie-lines, within 10 iterations. During the partitioning, the reference bus, bus 189, was assigned to Region 2 (mpc_regionR2). No buses were omitted, and the total number of buses across all regions matches the original MATPOWER dataset, with 200 buses, 245 branches, and 49 generators. The consistency of branch count, generator count, and overall system topology confirms the fidelity of the partitioning process. The final multi-region topology has been visualized and is shown in Fig. 2. The corresponding partitioned files are saved as pglib_opf_case200_tamu_3regions.mat and pglib_opf_case200_tamu_3regions.m. Both distributed DC OPF and AC OPF solvers have been executed on this partitioned data, and their convergence trends are illustrated through the worst primal residuals presented in Fig. 3 with $\rho=1000$.
Figure 2: Topology of pglib_opf_case200_tamu partitioned into three regions with 10 tie-lines.
(a) DC OPF
(b) AC OPF
Figure 3: Worst primal residual over iterations for distributed DC OPF and AC OPF on pglib_opf_case200_tamu.
IV-B Case: pglib_opf_case300_ieee
The pglib_opf_case300_ieee system is a widely used benchmark from the IEEE PES Power Grid Library for optimal power flow studies. This test case represents the IEEE 300-bus system, which was initially provided in the IEEE Common Data Format and converted into MATPOWER format on September 20, 2004, using the cdf2matp utility. The dataset was sourced from the University of Washington’s power systems test case archive. The system corresponds to typical operational conditions, with a base power of 100 MVA and includes detailed modeling of generation, loads, and transmission elements.
Using our graph-based spectral clustering approach, we have partitioned the pglib_opf_case300_ieee system into three electrically coherent regions. Before clustering, the original bus numbering in the MATPOWER file was reordered to ensure consistent indexing. The graph topology of the network was verified to be fully connected, allowing for a valid partitioning process. The algorithm identified the best clustering configuration in 10 iterations, resulting in 3 regions interconnected through 12 tie-lines. The system’s slack bus, originally defined at bus 257, was successfully assigned to Region 1 (denoted as mpc_regionR1). After partitioning, the integrity of the dataset was confirmed: all 300 buses, 411 branches, and 69 generators were fully included in the regional breakdown, with no data omitted. The final partitioned system has been saved as both a .mat file and a script-based .m file named pglib_opf_case300_ieee_3regions. A visual representation of the multi-regional topology is shown in Fig. 4. The distributed DC OPF and AC OPF solvers were successfully executed on this dataset, and the convergence behavior is demonstrated through the worst primal residual plots shown in Figs. 5(a) and 5(b) with $\rho=1000$.
Figure 4: Topology of pglib_opf_case300_ieee partitioned into three regions with 12 tie-lines.
(a) DC OPF
(b) AC OPF
Figure 5: Worst primal residual over iterations for distributed DC OPF and AC OPF on pglib_opf_case300_ieee.
IV-C Case: pglib_opf_case500_tamu
The pglib_opf_case500_tamu case is a synthetic 500-bus power system in the IEEE PES Power Grid Library for OPF studies. This case was developed as part of the ARPA-E GRID DATA research initiative and is geographically based in the northwestern region of South Carolina. It incorporates a transmission network operating at 345 kV and 138 kV voltage levels, constructed to supply power to a synthetic load distribution that mirrors the population density of the modeled region. The system does not represent any real grid infrastructure but is created using statistical modeling techniques described in [19], ensuring structural similarities to real transmission systems while maintaining confidentiality.
Using our graph-based spectral clustering framework, the pglib_opf_case500_tamu benchmark system was partitioned into five electrically coherent regions. Before clustering, the bus numbers from the original MATPOWER file were analyzed to preserve consistent indexing. The graph representation of the system was confirmed to be fully connected, ensuring the validity of the partitioning procedure. The spectral clustering algorithm converged in 10 iterations and identified an optimal structure of 5 regions interconnected through 16 tie-lines. The reference bus, originally defined as bus 17, was successfully assigned to Region 2 (denoted as mpc_regionR2). Post-partitioning checks confirmed that all 500 buses, 597 branches, and 90 generators were preserved without omission or duplication. The partitioned data has been saved in both .mat and script-based .m formats under the name pglib_opf_case500_tamu_5regions. A visual illustration of the multi-regional layout is presented in Fig. 6. Distributed DC OPF and AC OPF solvers were successfully executed for this case, and the convergence profiles are illustrated by the worst primal residual curves in Figs. 7(a) and 7(b) with $\rho=1000$.
Figure 6: Topology of pglib_opf_case500_tamu partitioned into five regions with 16 tie-lines.
(a) DC OPF
(b) AC OPF
Figure 7: Worst primal residual over iterations for distributed DC OPF and AC OPF on pglib_opf_case500_tamu.
IV-D Case: pglib_opf_case1354_pegase
The pglib_opf_case1354_pegase dataset is a detailed test case that emulates the size and complexity of the European high-voltage transmission network. It includes 1,354 buses, 260 generators, and 1,991 branches operating at 220 and 380 kV voltage levels. While the data is fictitious and does not reflect real-world infrastructure, it is highly suitable for validating power system optimization tools and techniques. This benchmark was developed as part of the Pan European Grid Advanced Simulation and State Estimation (PEGASE) project. It incorporates unique modeling features such as asymmetric shunt elements and generator aggregations.
The pglib_opf_case1354_pegase benchmark system was partitioned into eight electrically coherent regions using our graph-based spectral clustering method. The original bus indices from the MATPOWER file were reordered to ensure consistency during partitioning. The graph representation of the power network was verified to be fully connected, ensuring that a valid and complete partitioning could be achieved. The clustering algorithm successfully identified an optimal partitioning configuration within 10 iterations, resulting in 8 regions interconnected through 62 tie-lines. The system’s reference bus, which is located at bus 640 in the original data, was assigned to Region 1 (denoted as mpc_regionR1). The integrity of the data was verified after partitioning: all 1354 buses, 1991 branches, and 260 generators were accurately distributed across the regions with no missing components. The partitioned data files were saved in both .mat and .m formats under the name pglib_opf_case1354_pegase_8regions. A graphical representation of the resulting multi-regional topology is provided in Fig. 8. Both distributed DC OPF and AC OPF solvers were applied to this case study, with convergence behavior illustrated via worst primal residual plots in Figs. 9(a) and 9(b) with $\rho=1000$.
Figure 8: Topology of pglib_opf_case1354_pegase partitioned into eight regions with 62 tie-lines.
(a) DC OPF
(b) AC OPF
Figure 9: Worst primal residual over iterations for distributed DC OPF and AC OPF on pglib_opf_case1354_pegase.
IV-E Case: pglib_opf_case2383wp_k
The pglib_opf_case2383wp_k dataset models the Polish national transmission grid under peak load conditions during the winter of 1999–2000. It includes a comprehensive representation of 400, 220, and 110 kV networks and originates from a reduced version of the broader 7500+ bus European UCTE system. To simplify the model while preserving essential characteristics, tie lines connecting to neighboring countries were replaced by artificial load or generator buses. Generator aggregation and cost approximations were applied to reflect dispatchability in the Polish energy market.
Using the proposed graph-based spectral clustering method, the pglib_opf_case2383wp_k benchmark system was partitioned into twelve electrically meaningful regions. The underlying graph structure of the power system was verified to be fully connected, confirming the feasibility of multi-regional partitioning. The algorithm determined the optimal clustering within 10 iterations, producing 12 regions linked by 99 inter-regional tie-lines. The system’s slack bus, originally defined at bus 18 in the MATPOWER case, was assigned to Region 4 (referred to as mpc_regionR4).
Following partitioning, a full consistency check verified that all 2383 buses, 2896 branches, and 327 generators were preserved and correctly distributed across the regional datasets. No components were omitted during the process. The results were stored as both a .mat file and an executable .m script, saved under the name pglib_opf_case2383wp_k_12regions. The partitioned network topology is visually represented in Fig. 10. The distributed DC OPF and AC OPF solvers are implemented and their convergence behavior is shown through the worst primal residual plots in Figs. 11(a) and 11(b) with $\rho=2000$.
Figure 10: Topology of pglib_opf_case2383wp_k partitioned into twelve regions with 99 tie-lines.
(a) DC OPF
(b) AC OPF
Figure 11: Worst primal residual over iterations for distributed DC OPF and AC OPF on pglib_opf_case2383wp_k.
IV-F Case: pglib_opf_case3012wp_k
The pglib_opf_case3012wp_k dataset provides a detailed model of the Polish transmission grid during the winter 2007–2008 evening peak demand period. It represents the operational state of the 400, 220, and 110 kV networks and is a high-fidelity snapshot of the Polish system during that time. As part of the Power Grid Library benchmark suite, it is widely used for testing and validating optimization algorithms in realistic, large-scale scenarios.
Using the proposed graph-based spectral clustering framework, the pglib_opf_case3012wp_k benchmark system was partitioned into twelve electrically coherent regions. Before partitioning, the original MATPOWER bus numbers were reordered to enable consistent indexing throughout the graph representation. The full graph connectivity was verified to ensure that all buses could be properly included in the clustering process. Within 10 iterations, the partitioning algorithm identified an optimal configuration that divides the system into 12 regions connected by 100 inter-regional tie-lines. The slack bus, located initially at bus 37 in the MATPOWER data, was assigned to Region 12 (denoted as mpc_regionR12).
Following the partitioning process, a verification step confirmed that the distributed representation included all 3012 buses, 3572 branches, and 502 generators. These components were preserved without omission or duplication. The resulting data was saved both as a .mat file and as a script-based .m file under the name pglib_opf_case3012wp_k_12regions. A visual overview of the final partitioned topology is shown in Fig. 12. Using this distributed dataset, the DC OPF and AC OPF solvers were executed successfully, with their convergence behavior illustrated by the worst primal residual plots shown in Figs. 13(a) and 13(b) with $\rho=2000$.
Figure 12: Topology of pglib_opf_case3012wp_k partitioned into twelve regions with 100 tie-lines.
(a) DC OPF
(b) AC OPF
Figure 13: Worst primal residual over iterations for distributed DC OPF and AC OPF on pglib_opf_case3012wp_k.
IV-G Case: pglib_opf_case4661_sdet
The pglib_opf_case4661_sdet dataset is a large-scale synthetic power system model with 4,661 buses. It was created using the Sustainable Data Evolution Technology tool and is publicly available through the eGRID Data portal. This model is part of the PGLib OPF v19.05 benchmark suite and is provided in the public domain to support research in scalable and realistic power system analysis.
The pglib_opf_case4661_sdet benchmark system was partitioned into fourteen electrically coherent regions using the proposed graph-based spectral clustering methodology. Before partitioning, the original bus numbering in the MATPOWER data was reordered to enable consistent and systematic region assignment. The underlying graph representation of the network was verified to be fully connected, ensuring that all components were eligible for inclusion in the partitioning process. After 10 iterations, the algorithm identified an optimal clustering configuration with 14 regions and 284 inter-regional tie-lines. The reference (slack) bus, originally located at bus 2625 in the MATPOWER data, was assigned to Region 3 (denoted as mpc_regionR3).
A validation step was performed after partitioning to confirm the structural integrity of the distributed dataset. It verified that all 4661 buses, 5997 branches, and 1176 generators from the original MATPOWER case were successfully preserved and assigned across the generated regional files. The partitioned data has been saved in both .mat and script-based .m formats under the name pglib_opf_case4661_sdet_14regions. A visual summary of the multi-regional topology is presented in Fig. 14. Subsequently, the distributed DC OPF and AC OPF solvers were executed on the partitioned system, and their convergence characteristics are shown through the worst primal residual plots in Figs. 15(a) and 15(b) with $\rho=4000$.
Figure 14: Topology of pglib_opf_case4661_sdet partitioned into fourteen regions with 284 tie-lines.
(a) DC OPF
(b) AC OPF
Figure 15: Worst primal residual over iterations for distributed DC OPF and AC OPF on pglib_opf_case4661_sdet.
IV-H Case: pglib_opf_case6468_rte
The pglib_opf_case6468_rte [20] dataset provides a realistic snapshot of the French high voltage and very high voltage transmission network in 2013. It contains 6,468 buses, 1,295 generators, and 9,000 branches, offering a highly detailed and complex test case for power system optimization. While the data are based on real grid topologies, they are intended solely for academic and research purposes and should not be used for operational or planning decisions related to the French or European power grids. This dataset is part of the IEEE PES Power Grid Library (PGLib-OPF v19.05) and was developed by Réseau de Transport d’Électricité (RTE), the French transmission system operator. It supports validating optimization algorithms, mathematical modeling techniques, and simulation tools in large-scale power system analysis.
The pglib_opf_case6468_rte benchmark was successfully partitioned into electrically coherent regions using our spectral clustering-based partitioning algorithm. This method ensured that the network’s graph structure remained fully connected. The best clustering outcome was identified within 10 iterations, resulting in 16 regions interconnected by 204 inter-regional tie-lines. Bus renumbering was applied to the original MATPOWER case to enable consistent indexing across the partitioning workflow. The system’s slack bus, labeled as bus 4733, was assigned to Region 6 (mpc_regionR6).
All buses, branches, and generators from the original MATPOWER dataset were accounted for in the partitioned output, confirming the completeness and validity of the partitioning. Specifically, the resulting distributed dataset includes all 6468 buses, 9000 branches, and 1295 generators from the original case. The partitioned version has been saved in both .mat and script-based .m formats under the name pglib_opf_case6468_rte_16regions. A visualization of the region-level partitioned topology is presented in Fig. 16. To validate the partition, we applied the distributed DC OPF and AC OPF solvers, which successfully converged. Their worst primal residual plots over ADMM iterations are shown in Fig. 17 with $\rho=6000$.
Figure 16: Topology of pglib_opf_case6468_rte partitioned into 16 regions with 204 tie-lines.
(a) DC OPF
(b) AC OPF
Figure 17: Worst primal residual over iterations for distributed DC OPF and AC OPF on pglib_opf_case6468_rte.
IV-I Case: pglib_opf_case9241_pegase
The pglib_opf_case9241_pegase dataset provides a highly detailed representation of the European high voltage transmission network. It includes 9,241 buses, 1,445 generators, and 16,049 branches operating across multiple voltage levels, including 750, 400, 380, 330, 220, 154, 150, 120, and 110 kV. Although the dataset is fictitious and does not reflect real-world data, it is valuable for validating optimization algorithms and simulation tools in large-scale power system research. The data were released as part of the Pan European Grid Advanced Simulation and State Estimation (PEGASE) project [21].
The pglib_opf_case9241_pegase benchmark represents one of the largest systems in the MATPOWER repository and was successfully partitioned using our graph-based spectral clustering methodology. The partitioning process ensured complete inclusion of all 9241 buses and confirmed the full connectivity of the underlying power grid graph. The best clustering configuration was determined within 10 iterations, resulting in 18 regions connected by 197 inter-regional tie-lines. The system’s reference bus, numbered 4231, was assigned to Region 17 (mpc_regionR17). The reorganization of the data preserves the integrity of the original MATPOWER structure, with a total of 16049 branches and 1445 generators fully accounted for across the distributed regions. The partitioned results were saved in both .mat and .m formats under the name pglib_opf_case9241_pegase_18regions. A visual representation of the multi-region topology is provided in Fig. 18. This dataset has been validated by executing the distributed DC OPF and AC OPF solvers, with the convergence of the worst primal residuals depicted in Fig. 19 with $\rho=8000$.
Figure 18: Topology of pglib_opf_case9241_pegase partitioned into 18 regions with 197 tie-lines.
(a) DC OPF
(b) AC OPF
Figure 19: Worst primal residual over iterations for distributed DC OPF and AC OPF on pglib_opf_case9241_pegase.
V Conclusion
This work presented DPLib, a reproducible and open-source benchmark library tailored for distributed and decentralized power system studies. DPLib addresses the gap in standardized datasets for distributed OPF research by offering a graph-based partitioning toolkit and over 20 multi-region test cases derived from MATPOWER. The library is complemented by modular ADMM-based DC and AC OPF solvers, which enable scalable validation across diverse case sizes. Numerical results demonstrate the accuracy of the generated test cases and convergence of the solvers, making DPLib a practical foundation for future research in scalable, privacy-preserving, and resilient power system analysis.