### v3

# DPLib: A Standard Benchmark Library for Distributed Power System Analysis and Optimization

Abstract
DPLib is an open-source MATLAB-based benchmark library created to support research and development in distributed and decentralized power system analysis and optimization. Distributed and decentralized methods offer scalability, privacy preservation, and resilience to single points of failure, making them increasingly important for modern power systems. However, unlike centralized tools such as PGLib, no general-purpose, reproducible data library package currently exists for distributed power system studies. DPLib, available at GitHub, fills this gap by providing a standard power system library featuring over 20 multi-region benchmark test cases of varying sizes, along with a graph-based partitioning toolkit that decomposes any MATPOWER test system into multiple regions. The partitioning toolkit, an easy-to-use MATLAB code, generates standardized .mat and .m files, along with region visualizations for intuitive understanding. We also provide modular, easy-to-use distributed optimal power flow (OPF) solvers: an alternating direction method of multipliers (ADMM)-based DC OPF solver implemented in YALMIP, and an ADMM-based AC OPF solver leveraging IPOPT. These solvers validate the generated test systems for distributed optimization applications. Numerical results validate the generated test cases, establishing DPLib as a foundation for reproducible distributed power system research.
I Introduction
Centralized frameworks have long served as the dominant approach for modeling and solving large-scale power system problems, including optimal power flow (OPF), planning, control, and unit commitment. A key enabler behind the success and widespread adoption of centralized optimization and analysis in power systems is the availability of standardized benchmark test cases and libraries. Among these, MATPOWER has emerged as a widely accepted benchmark tool, offering consistent, well-structured data for centralized studies [22]. Its role in enabling reproducibility, fair performance comparisons, and methodological development has been fundamental to the evolution of centralized power system analysis.
For centralized studies, for instance, [11] proposed convex relaxations for AC OPF and validated their results on MATPOWER test cases. [16] provided a comprehensive review of relaxation techniques for OPF, with all benchmarks sourced from MATPOWER. [2] introduced QC relaxation for OPF using MATPOWER data. More recently, [19] developed robust optimization methods for corrective control under uncertainty, relying on MATPOWER cases for validation. [1] proposed the NESTA archive to enrich MATPOWER with more realistic data, yet retained full MATPOWER compatibility. These works and many more underscore the importance of a common, reliable data standard—facilitated by MATPOWER—that allows researchers to validate and compare centralized optimization algorithms under consistent and reproducible conditions.
In the last decade, distributed and decentralized computation has gained significant momentum in the power systems community, especially in applications such as distributed OPF, voltage control, state estimation, and demand-side management [15, 10, 6]. These include studies that apply alternating direction method of multipliers (ADMM) decomposition [10, 4], distributed voltage control [14], parallel and distributed OPF [21, 8], consensus-based optimization [18], and cooperative control frameworks [3, 7, 9]. Additional contributions include scalable distributed frameworks for transmission networks [12], graph-based coordination strategies [15], and distributed control in microgrids [5]. These methods offer inherent advantages in scalability, privacy preservation, and resilience to single points of failure.
However, as the literature on distributed techniques expands, a critical gap remains: the lack of unified, publicly available benchmark test cases tailored for distributed power system studies. Without such a standard library, researchers must often construct ad hoc datasets or apply arbitrary partitioning schemes, which hampers reproducibility and limits the comparability of different distributed algorithms. There is a strong need for a publicly available and community-accepted library of distributed test cases that allows researchers to validate, compare, and benchmark their distributed optimization or control methods under a common, well-defined framework.
In response to this need, we develop a distributed power system library, DPLib, that addresses the lack of standardized resources for distributed power system studies by providing a comprehensive library of over 20 benchmark test cases, ranging from 73 to 9241 buses. These multi-region systems are generated using the DPLib graph-based partitioning toolkit that captures the corresponding multi-regional network topology. Each case includes the necessary structural and connectivity data to support distributed analysis. All test cases have been verified using distributed OPF problems to demonstrate their practicality and confirm their suitability for reproducible and scalable distributed optimization.
In addition to a standard library, a key requirement is a general and user-friendly partitioning tool capable of processing any standard MATPOWER case and dividing it into subregions based on user-defined parameters, such as the number of regions. An effective tool should preserve the integrity of the original network, identify and document inter-regional tie-lines, and produce output in a structured format consistent with centralized MATPOWER cases. This functionality enables users to generate customized test cases while maintaining compatibility with a standard data structure. The DPLib graph-based partitioning toolkit fulfills all of these requirements.
While the distributed library includes several pre-partitioned and validated datasets, researchers may also wish to test their customized partitions generated by the DPLib graph-based toolkit. To support this, DPLib provides ADMM-based solvers for both DC OPF and AC OPF that are fully compatible with the partitioned data structure produced by the tool. These solvers take the partitioned datasets as input, iteratively solve the regional subproblems, enforce tie-line consistency through Lagrangian multipliers, and monitor convergence metrics, enabling reliable validation of user-defined partitions.
The main contribution of this work is the introduction of DPLib, a unified framework that establishes the first standard distributed data library for power system applications. Besides the library, DPLib consists of three core tools: a graph-based partitioning tool that decomposes any MATPOWER case into multiple regions using spectral clustering and k-means, enabling the creation of distributed datasets; a modular ADMM-based distributed DC OPF solver implemented in YALMIP [13]; and a distributed AC OPF solver using ADMM and IPOPT [20]. The DC OPF and AC OPF solvers are designed to validate and benchmark the partitioned datasets under realistic operating conditions. Together, these tools support reproducibility, scalability, and algorithm testing for distributed power system applications, addressing a key gap in current research infrastructure.
The remainder of the paper is organized as follows. Section II introduces the graph-based modeling of power networks and details the DPLib partitioning toolkit. Section III describes the ADMM-based DC and AC OPF solvers used to validate partitioned datasets. Section IV presents the benchmark case studies generated and tested using DPLib. Finally, Section V concludes the paper.
II Graph-Based Partitioning and Toolkit Description
In many power systems studies, realistic standard test cases in the MATPOWER format are commonly used for testing and validation. This practice ensures that developed methods are evaluated on realistic cases that reflect the complexity of actual power systems. To maintain this standard, researchers working on distributed algorithms should also rely on realistic standard test systems and datasets rather than arbitrarily constructed or manually partitioned grids.
In practice, real power networks often exhibit a naturally regional structure. For example, a network covering several regions will typically have strong internal connectivity within each region and only a few tie-lines connecting areas to one another. This results in regions that are strongly connected internally but weakly connected to the other areas. A proper partitioning method should be able to observe these loosely connected boundaries. We use graph-based partitioning to provide a principled and effective framework for capturing such structures and dividing the network in a way that aligns with real systems’ physical and topological characteristics.
II-A Power Network Modeling and Spectral Clustering
To prepare the network for spectral partitioning, the power system is modeled as an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where buses form the nodes and transmission lines form the edges. The graph representation is encoded through the adjacency matrix $A$, and two forms of this matrix may be constructed depending on how connectivity is interpreted.
In its simplest form, the network is viewed purely through its topological structure. Each transmission line indicates the presence of a connection, and the adjacency matrix is as
$A_{ij}=\begin{cases}1,&\text{if a line connects buses }i\text{ and }j,\\ 0,&\text{otherwise}.\end{cases}$ (1)
This unweighted representation treats all lines uniformly and emphasizes the combinatorial shape of the network. Such a formulation is robust to data irregularities, reflects the way regional boundaries appear in practice when tie-lines are sparse, and preserves the graph-theoretic properties that spectral clustering exploits most directly.
A second option introduces electrical heterogeneity explicitly. In this case, the weight of an edge is determined by parameters such as line reactance $x_{ij}$, line charging susceptance $b_{ij}$, and thermal rating $S_{\max,ij}$. After appropriate normalization, a composite quantity is assigned to each existing edge,
$A_{ij}=w(x_{ij},b_{ij},S_{\max,ij}),$ (2)
so the graph captures not only the presence of a connection but also its relative electrical strength. This weighted representation highlights tightly coupled corridors and down-weights weak or limited-capacity lines, and therefore encodes additional physical detail into the graph used for clustering.
Although the toolkit supports a weighted Laplacian that incorporates electrical heterogeneity, we treat this option as an auxiliary feature rather than the default mode. Weighted formulations can highlight electrically strong corridors, but doing so requires a choice of normalization, scaling, and aggregation of heterogeneous parameters—decisions that are not standardized across power-system studies. Because the purpose of DPLib is to produce baseline multi-region datasets for benchmarking distributed algorithms, we use the unweighted formulation to ensure that all results are comparable, repeatable, and free of case-dependent parameter tuning. Researchers who wish to emphasize electrical strength rather than topological connectivity may enable the weighted variant, which the toolkit provides for completeness.
Both constructions lead to the degree matrix $D$ defined by $D_{ii}=\sum_{j}A_{ij}$, and the Laplacian L=DA.L=D-A. The Laplacian matrix $L$ encodes the power network connectivity and topology. For a connected graph, $L$ is symmetric and positive semi-definite, meaning all its eigenvalues are non-negative. Importantly, the smallest eigenvalue of $L$ is always zero, corresponding to the trivial uniform eigenvector.
To divide the system into $k$ regions, we use spectral clustering, which leverages the properties of $L$. The first step is to compute the eigenvalues and eigenvectors of $L$. Among these, the $k$ smallest nonzero eigenvalues are selected. Each eigenvalue is associated with an eigenvector that varies smoothly across the graph, meaning that buses that are strongly connected in the physical system will have similar eigenvalues. Therefore, these eigenvectors reflect the underlying structure of the network.
We form a matrix $U\in\mathbb{R}^{n\times k}$ by placing the $k$ selected eigenvectors as columns. Each row $u_{i}\in\mathbb{R}^{k}$ corresponds to a bus and represents it in a new $k$-dimensional spectral space, where geometrical proximity reflects topological connectivity.
To assign buses into one of $k$ regions, we apply the k-means clustering algorithm to the rows of $U$. This unsupervised learning algorithm operates by:
  • Randomly initializing $k$ centroids (in our case, from the rows of $U$);
  • Assigning each bus to the nearest centroid based on Euclidean distance;
  • Recalculating the centroid of each cluster;
  • Repeating the assignment and centroid update steps until convergence.
The clustering result yields a partition of all buses into $k$ regions. Each cluster groups strongly connected buses, leading to balanced regions where each has its generation. After forming the clusters, inter-regional tie-lines (lines that connect buses belonging to different clusters) are identified. To improve quality, the algorithm repeats the clustering multiple times and select the configuration that minimizes the number of inter-regional tie-lines.
II-B DPLib Partitioning Toolkit
The DPLib partitioning toolkit implements the graph-based clustering framework described in the previous subsection and automates the full workflow for constructing multi-region MATPOWER datasets. The toolkit supports both adjacency representations introduced earlier, allowing the Laplacian to be formed either from the purely topological structure of the network or from a weighted graph that incorporates electrical parameters. In the unweighted mode, each transmission line contributes a unit entry to the adjacency matrix, so the Laplacian reflects the combinatorial connectivity of the system. In the weighted mode, each existing edge is assigned a composite weight constructed from line reactance, line charging susceptance, and thermal capacity after normalization, allowing the Laplacian to embed relative electrical strength directly into the spectral geometry.
Although both modes are supported, the unweighted Laplacian is used as the default for generating standardized multi-region test cases. This choice ensures that the resulting partitions depend only on the physical topology of the MATPOWER system and remain independent of parameter scaling, model variations, or dataset-dependent numerical ranges. The unweighted construction also aligns with how regional boundaries appear in real-world transmission networks, where geographical separations and sparse tie-lines primarily determine structural decomposition. Using the unweighted version therefore provides consistent, reproducible partitions across different datasets and preserves compatibility with studies that compare methods on topologically equivalent cases.
The workflow proceeds as follows. The MATPOWER case is loaded, and if the original bus numbering is not sequential, all buses, branches, and generators are renumbered from $1$ to $n$. Depending on the selected mode, the adjacency matrix is constructed using either the binary definition or the weighted formulation. The degree matrix and Laplacian are then formed, and the $k$ Laplacian eigenvectors associated with the smallest nonzero eigenvalues are computed to obtain the spectral embedding. Multiple runs of $k$-means are applied to the rows of this matrix to obtain a stable clustering, and the configuration with the smallest number of inter-regional tie-lines is selected. For the chosen partition, regional subcases are created by extracting internal buses, branches, generators, and cost data, and each region is locally renumbered to maintain consistency. Tie-lines connecting different regions are identified and stored together with their full MATPOWER branch data.
The completed partition is exported in several formats. A .mat file contains all regional subcases and a table describing every inter-regional tie-line. A corresponding .m script reproduces the same structures in a clear and readable format. A graphical visualization is produced in .png format, showing regions arranged in a force-directed layout together with annotated tie-lines. These outputs provide a consistent, reproducible, and easily verifiable multi-region dataset that can be used directly in distributed studies. The algorithm of the partitioning toolkit, illustrating the complete data processing pipeline, is provided in Algorithm 1.
Algorithm 1 DPLib Partitioning Toolkit
1: Input MATPOWER case name and desired number of regions $k$
2: Load MATPOWER data; if bus numbering is non-sequential, renumber buses and update references
3: Construct adjacency matrix AA:
  1. 1.
    Unweighted: $A_{ij}=1$ if a branch connects $i$ and $j$
  2. 2.
    Weighted: assign $A_{ij}=w(x_{ij},b_{ij},S_{\max,ij})$
4: Form degree matrix $D$ and Laplacian L=DAL=D-A
5: Compute the Laplacian eigenvectors associated with the $k$ smallest nonzero eigenvalues and form $U$
6: For multiple attempts: apply $k$-means to rows of $U$, build candidate regions, identify tie-lines, and discard candidates without generators
7: Select the valid partition with the minimum number of tie-lines
8:for each region do
9:  Extract internal buses, branches, generators, and cost data; locally renumber buses
10:end for
11: Identify the region of the slack bus and record all inter-regional tie-lines
12: Export regional cases to .mat and .m files and generate topology visualization
III Validating DPLib Test Systems Using OPF
To ensure that the generated multi-region test cases are suitable for distributed power system studies, it is essential to validate them through power system analysis. Although the DPLib partitioning toolkit generates distributed variants of MATPOWER test systems applicable to a wide range of power system studies, including control, state estimation, or stability analysis, this section evaluates these test systems for distributed OPF applications only. OPF has been selected as a representative benchmarking problem because of its maturity, availability of centralized references, and compatibility with distributed solvers due to over a decade of research in this area. We have developed general ADMM-based solvers for DC and AC OPF problems that operate directly on the test cases produced by the DPLib partitioning toolkit. These solvers are designed to validate the feasibility, consistency, and convergence of partitioned datasets, demonstrating the suitability of these test systems. Nevertheless, users may use alternative solvers or utilize these provided tools to verify and explore their own partitioned datasets.
III-A Distributed OPF Formulation
We formulate the OPF problem in both DC and AC settings for each region $r$. Each region solves a local OPF problem while maintaining consistency at the boundaries via duplicated (virtual) variables and tie-line constraints [8]. These virtual variables (buses) enable each region to model tie-line interactions locally without requiring direct enforcement of global constraints. The overall structure is illustrated in Fig. 1(a), where a two-region power grid is shown with inter-regional tie-lines in red. Fig. 1(b) further illustrates how virtual buses (shown in red) are added to represent shared variables.
(a) Grid with two regions and red lines are inter-regional transmission tie-lines
(b) Decomposition into two regions with virtual buses shown in red
Figure 1: Illustration of a partitioned power system
III-A1 DC OPF Formulation
The DC OPF for region $r$ is [8]:
$\min_{\boldsymbol{x}^{r}}\sum_{i\in\Omega^{r}_{g}}f(p^{g}_{i})$ (3a)
s.t.
$\qquad\quad\underline{p}_{i}^{g}\leqslant p^{g}_{i}\leqslant\bar{p}_{i}^{g},\quad\forall i\in\Omega^{r}_{g}$ (3b)
$\qquad\quad p^{g}_{i}-p^{d}_{i}=\sum_{j\in\Omega^{r}_{i}}\frac{\theta_{i}-\theta_{j}}{x_{ij}}+\sum_{j\in\Omega^{r,\textrm{in}}_{i}}\frac{\theta_{i}-\tilde{\theta}_{j}}{x_{ij}},\quad\forall i\in\Omega^{r}_{b}$ (3c)
$\qquad\quad\underline{p}_{ij}\leqslant\frac{\theta_{i}-\theta_{j}}{x_{ij}}\leqslant\bar{p}_{ij},\quad\forall ij\in\Omega^{r}_{l}$ (3d)
$\qquad\quad\underline{p}_{ij}\leqslant\frac{\theta_{i}-\tilde{\theta}_{j}}{x_{ij}}\leqslant\bar{p}_{ij},\quad\forall ij\in\Omega^{r,\textrm{in}}_{l}$ (3e)
$\qquad\quad\theta_{\text{ref }}=0,$ (3f)
$\qquad\quad\tilde{\theta}_{j}=\theta_{j},\quad\forall j:ij\in\Omega^{r,\textrm{in}}_{l}$ (3g)
$\qquad\quad\theta_{i}=\tilde{\theta}_{i},\quad\forall i:ij\in\Omega^{r,\textrm{in}}_{l}$ (3h)
The decision vector for region r contains the active power generations $p^{g}_{i}$ for all generating units $i\in\Omega^{r}_{g}$ and the bus voltage phase angles $\theta_{i}$ for all buses $i\in\Omega^{r}_{b}$, giving $x^{r}\in\mathbb{R}^{n^{r}_{g}+n^{r}_{b}}$. The objective in (3a) minimizes the local generation cost $f(p^{g}_{i})$. Constraint (3b) applies the minimum and maximum generation limits for all units in the set $\Omega^{r}_{g}$. The nodal balance equation (3c) uses the demand $p^{d}_{i}$, the reactance $x_{ij}$ of each line, the set of internal neighboring buses $\Omega^{r}_{i}$, and the set of inter-regional neighboring buses $\Omega^{r,\mathrm{in}}_{i}$, and incorporates virtual boundary angles $\tilde{\theta}_{j}$ that represent buses belonging to adjacent regions. Constraints (3d)–(3e) impose the power-flow limits $\underline{p}_{ij}$ and $\bar{p}_{ij}$ on all internal lines in $\Omega^{r}_{l}$ and all tie-lines in $\Omega^{r,\mathrm{in}}_{l}$. Constraint (3f) selects the regional reference bus. Finally, constraints (3g)–(3h) enforce angle consistency across inter-regional connections by matching the real boundary-bus angles with their corresponding virtual angles.
III-A2 AC OPF Formulation
The AC OPF for region $r$ [8]:
$\min_{\boldsymbol{x}^{r}}\sum_{i\in\Omega^{r}_{g}}f(p^{g}_{i})$ (4a)
s.t.
$\qquad\underline{p}_{i}^{g}\leqslant p^{g}_{i}\leqslant\bar{p}_{i}^{g},\quad\forall i\in\Omega^{r}_{g}$ (4b)
$\qquad\underline{q}_{i}^{g}\leqslant q^{g}_{i}\leqslant\bar{q}_{i}^{g},\quad\forall i\in\Omega^{r}_{g}$ (4c)
$\qquad p^{g}_{i}-p^{d}_{i}=\sum_{j\in\Omega^{r}_{i}}\left(p_{ij,f}+p_{ij,t}\right)+\sum_{j\in\Omega^{r,\text{in}}_{i}}{p}_{ij,\xi},\quad\forall i\in\Omega^{r}_{b}$ (4d)
$\qquad q^{g}_{i}-q^{d}_{i}=\sum_{j\in\Omega^{r}_{i}}\left(q_{ij,f}+q_{ij,t}\right)+\sum_{j\in\Omega^{r,\text{in}}_{i}}{q}_{ij,\xi},\quad\forall i\in\Omega^{r}_{b}$ (4e)
$\qquad p_{ij,f}=G_{ij,1}v_{i}^{2}+G_{ij,2}v_{i}v_{j}\cos(\theta_{i}-\theta_{j})$
$\qquad\qquad\quad+B_{ij,2}v_{i}v_{j}\sin(\theta_{i}-\theta_{j}),\quad\forall ij\in\Omega^{r}_{l}$ (4f)
$\qquad q_{ij,f}=-B_{ij,1}v_{i}^{2}-B_{ij,2}v_{i}v_{j}\cos(\theta_{i}-\theta_{j})$
$\qquad\qquad\quad+G_{ij,2}v_{i}v_{j}\sin(\theta_{i}-\theta_{j}),\quad\forall ij\in\Omega^{r}_{l}$ (4g)
$\qquad p_{ij,t}=G_{ij,4}v_{j}^{2}+G_{ij,3}v_{i}v_{j}\cos(\theta_{j}-\theta_{i})$
$\qquad\qquad\quad+B_{ij,3}v_{i}v_{j}\sin(\theta_{j}-\theta_{i}),\quad\forall ij\in\Omega^{r}_{l}$ (4h)
$\qquad q_{ij,t}=-B_{ij,4}v_{j}^{2}-B_{ij,3}v_{i}v_{j}\cos(\theta_{j}-\theta_{i})$
$\qquad\qquad\quad+G_{ij,3}v_{i}v_{j}\sin(\theta_{j}-\theta_{i}),\quad\forall ij\in\Omega^{r}_{l}$ (4i)
$\qquad p_{ij,f}^{2}+q_{ij,f}^{2}\leq\bar{p}_{ij}^{2}+\bar{q}_{ij}^{2},\quad\forall ij\in\Omega^{r}_{l}$ (4j)
$\qquad p_{ij,t}^{2}+q_{ij,t}^{2}\leq\bar{p}_{ij}^{2}+\bar{q}_{ij}^{2},\quad\forall ij\in\Omega^{r}_{l}$ (4k)
$\qquad{p}^{2}_{ij,\xi}+{q}^{2}_{ij,\xi}\leq\bar{p}_{ij}^{2}+\bar{q}_{ij}^{2},\quad\forall ij\in\Omega^{r,\textrm{in}}_{l}$ (4l)
$\qquad\underline{\Delta\theta}_{ij}\leq\theta_{i}-\theta_{j}\leq\overline{\Delta\theta}_{ij},\quad\forall ij\in\Omega^{r}_{l}$ (4m)
$\qquad\underline{\Delta\theta}_{ij}\leq\theta_{i}-\tilde{\theta}_{j}\leq\overline{\Delta\theta}_{ij},\quad\forall ij\in\Omega^{r,\textrm{in}}_{l}$ (4n)
$\qquad\theta_{\text{ref}}=0,$ (4o)
$\qquad\tilde{\theta}_{j}=\theta_{j},\quad\forall j:ij\in\Omega^{r,\textrm{in}}_{l}$ (4p)
$\qquad\theta_{i}=\tilde{\theta}_{i},\quad\forall i:ij\in\Omega^{r,\textrm{in}}_{l}$ (4q)
$\qquad\tilde{v}_{j}=v_{j},\quad\forall j:ij\in\Omega^{r,\textrm{in}}_{l}$ (4r)
$\qquad v_{i}=\tilde{v}_{i},\quad\forall i:ij\in\Omega^{r,\textrm{in}}_{l}$ (4s)
The AC OPF subproblem for region r uses the same generation set $\Omega^{r}_{g}$ and bus set $\Omega^{r}_{b}$, where the decision variables include active and reactive generations $p^{g}_{i}$ and $q^{g}_{i}$, bus voltage magnitudes $v_{i}$, and phase angles $\theta_{i}$, together with virtual boundary variables $\tilde{\theta}_{j}$ and $\tilde{v}_{j}$ for all inter-regional neighbors in $\Omega^{r,\mathrm{in}}_{i}$. The objective in (4a) minimizes the generation cost over all units in $\Omega^{r}_{g}$. Generator active and reactive limits are enforced in (4b)–(4c). Power-balance constraints (4d)–(4e) use local demands $p^{d}_{i}$ and $q^{d}_{i}$, as well as flows from internal neighbors $j\in\Omega^{r}_{i}$ and inter-regional neighbors $j\in\Omega^{r,\mathrm{in}}_{i}$, where $p_{ij,f},p_{ij,t},q_{ij,f},q_{ij,t}$ represent the AC branch flows defined by the conductance and susceptance parameters $G_{ij,k}$ and $B_{ij,k}$ in (4f)–(4i). Thermal limits for internal lines $\Omega^{r}_{l}$ and tie-lines $\Omega^{r,\mathrm{in}}_{l}$ are enforced in (4j)–(4l). The phase-angle difference bounds in (4m)–(4n) use the parameters $\underline{\Delta\theta}_{ij}$ and $\overline{\Delta\theta}_{ij}$, and (4o) sets the reference bus angle. Finally, constraints (4p)–(4s) ensure consistency between the real and virtual boundary-bus variables, matching $\theta_{i}$ with $\tilde{\theta}_{i}$ and $v_{i}$ with $\tilde{v}_{i}$ across all inter-regional connections in $\Omega^{r,\mathrm{in}}_{l}$.
III-A3 ADMM Decomposition
The distributed OPF solver uses the sequential form of ADMM, where regions are updated one after another in a Gauss–Seidel manner. Each region solves its local OPF problem using the most recent boundary voltage information from neighbors that have already been updated in the current iteration, while using the previous-iteration values for neighbors that have not yet been updated. The shared variables are the voltage phase angles and voltage magnitudes of boundary buses. Their virtual copies represent the neighboring regions’ views of these boundary variables.
Before the ADMM iterations start, we also compute problem-dependent scaling factors for the boundary mismatches. For the phase angles, a characteristic scale $\theta_{\text{scale}}$ is obtained from the line reactances and thermal limits as
$\theta_{\max,\ell}\approx|x_{\ell}|\,\frac{\text{rateA}_{\ell}}{\text{baseMVA}},$
and $\theta_{\text{scale}}$ is chosen as a robust statistic (for example, the median) of $\{\theta_{\max,\ell}\}$ over all lines. For AC OPF, an additional voltage-magnitude scale $v_{\text{scale}}$ is defined from the bus voltage limits, for example as half the spread between the minimum and maximum allowed magnitudes. These scales are used to normalize the consensus residuals and make them dimensionless.
Consistency across regions is enforced by an augmented Lagrangian of the form
$L_{\rho}(\theta,\tilde{\theta},v,\tilde{v},\lambda_{\theta},\lambda_{v})=$ $f(p^{g})+\lambda_{\theta}^{\top}(\frac{\theta-\tilde{\theta}}{\theta_{\text{scale}}})+\lambda_{v}^{\top}(\frac{v-\tilde{v}}{v_{\text{scale}}})$
$+\tfrac{\rho_{\theta}}{2}\|\frac{\theta-\tilde{\theta}}{\theta_{\text{scale}}}\|^{2}+\tfrac{\rho_{v}}{2}\|\frac{v-\tilde{v}}{v_{\text{scale}}}\|^{2},$ (5)
[…]
During iteration $k$, region $r$ performs the primal update by solving its local AC or DC OPF subproblem. The virtual variables from neighbors are fixed parameters in this optimization. After the OPF solve, region $r$ updates its real boundary voltages $(\theta_{r}^{k+1},v_{r}^{k+1})$ and refreshes its virtual copies $(\tilde{\theta}_{r}^{k+1},\tilde{v}_{r}^{k+1})$ using the newest available values from neighboring regions. This constitutes the primal update step of ADMM. Once all regions finish their primal updates, the dual variables are updated based on normalized residuals. For each tie-line boundary bus, the updates are
$\lambda_{\theta}^{k+1}$ $=\lambda_{\theta}^{k}+\rho_{\theta}\frac{\theta^{k+1}-\tilde{\theta}^{k+1}}{\theta_{\text{scale}}},$ (6)
$\lambda_{v}^{k+1}$ $=\lambda_{v}^{k}+\rho_{v}\frac{v^{k+1}-\tilde{v}^{k+1}}{v_{\text{scale}}}.$ (7)
In the DC-OPF case only the angle channel is present, so the update for $\lambda_{v}$ and the associated terms are omitted.
Convergence is assessed using the worst-case normalized primal residual across all tie-lines. The primal residual at iteration $k+1$ is computed as
$r_{\text{pri}}^{k+1}=\max_{(i,j)\in\mathcal{E}_{\text{tie}}}\left\{\frac{|\theta_{i}^{k+1}-\tilde{\theta}_{i}^{k+1}|}{\theta_{\text{scale}}},\,\frac{|v_{i}^{k+1}-\tilde{v}_{i}^{k+1}|}{v_{\text{scale}}}\right\},$
matching the residual definition used in [8], up to the normalization by $(\theta_{\text{scale}},V_{\text{scale}})$. In the DC-OPF case only the angle term is used in the maximum. The ADMM iterations terminate once this residual falls below a prescribed tolerance.
To improve robustness across different network topologies and operating points, the penalty parameters are updated adaptively based on the relative sizes of the primal and dual residuals, following the heuristic in [17]. Let $r_{\text{pri}}^{k}$ and $r_{\text{dual}}^{k}$ denote the normalized primal and dual residuals at iteration $k$. For the distributed DC-OPF solver, a single penalty parameter is updated according to
$\rho^{k+1}=\begin{cases}\tau_{\text{incr}}\,\rho^{k},&\text{if }r_{\text{pri}}^{k}>\mu\,r_{\text{dual}}^{k},\\[1.99997pt] \rho^{k}/\tau_{\text{decr}},&\text{if }r_{\text{dual}}^{k}>\mu\,r_{\text{pri}}^{k},\\[1.99997pt] \rho^{k},&\text{otherwise},\end{cases}$
with chosen parameters $\mu$, $\tau_{\text{incr}}$, and $\tau_{\text{decr}}$. For the distributed AC-OPF solver, the same heuristic is applied separately to the angle and magnitude channels, i.e., to $\rho_{\theta}$ and $\rho_{v}$, using different balance factors for the $\theta$ and $v$ residuals.
To prevent numerical blow-up and avoid instability caused by excessively large penalties, we enforce an upper bound on the adaptive updates. Specifically, the penalty updates are terminated once the penalty reaches a prescribed maximum value: $\rho_{\mathrm{DC}}^{\max}$ for the distributed DC-OPF solver and $\rho_{\mathrm{AC}}^{\max}$ for the distributed AC-OPF solver. In practice, if an update would increase the current penalty beyond the corresponding cap, we set $\rho\leftarrow\rho_{\mathrm{DC}}^{\max}$ (DC) or $\rho_{\theta},\rho_{v}\leftarrow\rho_{\mathrm{AC}}^{\max}$ (AC) and disable further penalty adaptation thereafter. This safeguard ensures the algorithm does not diverge due to runaway growth of $\rho$.
Moreover, to avoid overly frequent parameter changes—which can introduce oscillations and hinder convergence—we apply a cooldown period after each penalty update. Concretely, once $\rho$ (DC) or $\rho_{\theta},\rho_{v}$ (AC) is updated, we postpone any subsequent residual-balance checks and penalty updates for the next $N_{\mathrm{wait}}$ iterations, where $N_{\mathrm{wait}}=20$ for cases with fewer than $2000$ buses and $N_{\mathrm{wait}}=40$ for cases with $2000$ buses or more. This heuristic empirically stabilizes the convergence behavior by allowing the primal and dual iterates to settle under the new penalty before further adaptation is attempted.
Overall, this adaptive strategy automatically increases the penalty when the primal residual dominates and decreases it when the dual residual dominates, promoting a better balance between consensus enforcement and dual convergence while remaining numerically stable through the maximum-penalty cap and the update cooldown mechanism.
Figure 2 summarizes the sequential ADMM workflow, highlighting the order of local OPF solves, voltage exchange, dual updates (with scaled and adaptively tuned penalties), and residual checks.
Initialize ADMMFor $r=1,\dots,K$: solve regional OPF and update boundary voltagesUpdate dual variables on tie-lines (scaled residuals, adaptive $\rho$)Compute worst-case normalized primal residualConverged?StopNext ADMM iterationyesno
Figure 2: Sequential ADMM loop for distributed OPF with scaled residuals and adaptive penalties
III-B General ADMM-Based OPF Solvers
We have developed two open-source general-purpose solvers for distributed OPF problems, one for DC OPF and another for AC OPF. Both solvers adopt ADMM to coordinate regional solutions and ensure consistency across tie-lines.
III-B1 DPLib DC OPF Solver
The DPLib DC OPF module implements a distributed DC OPF solver based on the ADMM. The MATLAB implementation, available in the file DistDCOPF.m and publicly accessible on GitHub, is designed for partitioned MATPOWER cases and supports the YALMIP optimization backends in MATLAB. The solver requires MATLAB R2020a or newer with MATPOWER installed; YALMIP must also be added to the MATLAB path.
At execution time, the user provides the partitioned MATPOWER file, the ADMM penalty parameter rho, a convergence tolerance, a maximum number of iterations, a $\rho_{\mathrm{DC}}^{\max}$, and optionally the centralized OPF cost for computing the optimality gap. YALMIP emphasizes modeling clarity and solver flexibility, and is suitable for rapid prototyping and debugging.
After receiving the user inputs, DistDCOPF.m loads all regional data, reconstructs inter-regional tie-lines, and augments each region’s branch matrix accordingly. Bus-angle variables, tie-line indices, and interface bus mappings are then initialized, along with the ADMM dual variables. During each ADMM iteration, every region solves its local DC OPF problem independently using the selected backend. The local objective minimizes the generation cost subject to regional constraints and tie-line consistency penalties. Following all local solves, the solver computes the angle mismatches on shared tie-lines, updates the dual variables, records the maximum primal residual, and checks the convergence condition. If a centralized optimal cost is available, the optimality gap is also computed and plotted. The procedure continues until convergence is reached or the iteration limit is met.
The solver supports any number of regions and enables reproducible distributed OPF studies. Its overall workflow is summarized in Algorithm 2.
Algorithm 2 DPLib Distributed OPF Solver
1: Input partitioned MATPOWER data, ADMM parameters rho, tolerance epsilon, maximum iterations
2: Load all regional data and global tie-line table
3: Augment each region with its corresponding inter-regional tie-line rows
4: Initialize regional variables, interface indices, and ADMM dual variables
5: Prepare regional consistency constraints and interface bus mappings
6:for each ADMM iteration do
7:  for each region do
8:   Solve regional OPF
9:  end for
10:  Compute tie-line angle mismatches across all region boundaries
11:  Update dual variables based on consistency violations
12:  Evaluate primal residual and check convergence criterion
13:  if convergence achieved then
14:   Break
15:  end if
16:end for
17: Visualize residuals, optimality gap (if applicable), and final regional solutions
III-B2 DPLib AC OPF Solver
The DPLib AC OPF solver provides a MATLAB implementation of a distributed ADMM-based AC OPF algorithm that operates on any power system previously partitioned using the DPLib framework. The solver is implemented in the publicly available file DistACOPF.m, accessible through the GitHub repository at GitHub. The code requires MATLAB R2020a or later, along with a properly configured installation of MATPOWER and IPOPT.
The function begins by requesting the partitioned .mat file and the user-defined ADMM parameters, including the penalty value $\rho$, the convergence tolerance, the maximum number of iterations, the $\rho_{\mathrm{AC}}^{\max}$, and optionally the centralized AC OPF cost for computing the optimality gap. After these inputs are provided, the solver loads the partitioned data and augments the branch matrices of each region with their associated tie-lines. It then initializes regional variables describing bus counts, voltage magnitudes and angles, and other quantities needed to enable the local AC OPF solves and the coordination steps.
The main ADMM loop is implemented within the DistACOPF.m function. In each iteration, every region independently solves its local nonlinear AC OPF using the helper function acopf, which is based on the formulation in equation (4). The updated voltage magnitudes and phase angles from each region are then exchanged virtually through the consistency constraints applied on the tie-lines. The dual variables are modified at each step using the mismatch in the angle and magnitude variables across the interfaces, ensuring that neighboring regions gradually reach agreement. The loop continues until the maximum primal residual across all tie-lines drops below the specified tolerance. If a centralized AC OPF cost is provided, the function also evaluates the optimality gap at every iteration to quantify the progress toward the global optimum.
Once the iterations terminate, the solver prints summary information and generates plots that illustrate the convergence behavior of the primal residual and, when applicable, the optimality gap. The complete algorithmic flow implemented in DistACOPF.m is summarized in Algorithm 2.
IV DPLib: Benchmark Test Cases
TABLE I: Partitioning summary with centralized DC/AC OPF costs.
System Size Partition Penalty Param. Centralized OPF Cost
Case Buses Branches Gens Regions Tie-lines Slack $\rho$ $\rho_{\mathrm{DC}}^{\max}$ $\rho_{\mathrm{AC}}^{\max}$ DC AC
pglib_opf_case200_tamu $200$ $245$ $49$ 33 $10$ R2 11 1e81108 1e81108 $34,653$ $27,558$
pglib_opf_case300_ieee $300$ $411$ $69$ 33 $12$ R1 11 1e81108 1e81108 $517,310$ $565,220$
pglib_opf_case500_tamu $500$ $597$ $90$ 55 $16$ R2 11 1e81108 1e81108 $91,841$ $72,578$
pglib_opf_case1354_pegase $1,354$ $1,991$ $260$ 88 $62$ R1 11 1e81108 1e81108 $1,218,028$ $1,258,844$
pglib_opf_case2869_pegase $2,869$ $4,582$ $510$ $12$ $95$ R6 11 1e51105 1e81108 $2,386,170$ $2,462,790$
pglib_opf_case4661_sdet $4,661$ $5,997$ $1,176$ $14$ $284$ R3 11 1e61106 1e61106 $2,217,130$ $2,255,540$
pglib_opf_case9241_pegase $9,241$ $16,049$ $1,445$ $18$ $197$ R17 11 1e61106 1e51105 $6,044,174$ $6,243,090$
This section presents a detailed description and verification of several DPLib benchmark systems and their corresponding multi-region topologies. All test cases originate from the PGLib-OPF library and are partitioned using the spectral clustering procedure of Algorithm 1 based on the unweighted Laplacian matrix. Before partitioning, every system undergoes bus renumbering when required, connectivity verification, slack-region assignment, and integrity checks to ensure that all buses, branches, and generators are preserved. The resulting number of regions, tie-lines, and solver parameters are summarized in Table I. All region-labeled MATPOWER files, verification codes, and the topology and verification figures included here are available in the DPLib GitHub repository.
To make the library easy to use and to avoid case-by-case tuning, the ADMM penalty parameters are initialized to $\rho^{0}=1$ for all benchmarks in both the distributed DC-OPF and AC-OPF solvers, after which $\rho$ is adjusted automatically via the adaptive residual-balancing rule where $\tau_{\mathrm{incr}}=2,\tau_{\mathrm{decr}}=0.5$. In the distributed DC-OPF solver, the adaptive update parameters are set internally to
$\mu=\max\!\left(1,\ \frac{15}{(n_{\mathrm{reg}}^{2})\cdot 0.2}\right),$ (8)
where $n_{\mathrm{reg}}$ denotes the number of regions. For the distributed AC-OPF solver, the same heuristic is applied separately to the angle and voltage-magnitude channels using
$\mu_{\theta}=\max\!\left(1,\ \frac{15}{(n_{\mathrm{reg}}^{2})\cdot 0.2}\right),\;\mu_{v}=\max\!\left(1,\ \frac{10}{(n_{\mathrm{reg}}^{2})\cdot 0.2}\right).$ (9)
These choices scale the residual-balance thresholds with the partition granularity while keeping fixed and interpretable increase/decrease factors, allowing the adaptive-$\rho$ mechanism to operate robustly across different network sizes and region counts with minimal user intervention.
The verification procedure is implemented in a MATLAB script that processes partitioned cases in a unified workflow. For each system, the script first loads the corresponding MATPOWER case and calls the centralized solvers run_dcopf_centralized and run_acopf_centralized. These functions compute reference DC OPF and AC OPF objective values and return solver status information. The centralized costs serve as benchmarking values to compute the optimality gap of the distributed solutions if the centralized DC OPF is feasible. The distributed DC OPF solver is run using the function DistDCOPF with prescribed tolerance and maximum iteration count. Similarly, the distributed AC OPF solver DistACOPF is called following the structure of Algorithm 2. For each distributed run, the script tracks the worst primal residual across tie-lines and the optimality gap with respect to the centralized objective value. Table I reports placeholder centralized DC and AC objective values for all systems. In the numerical experiments, for all cases where the centralized solvers return feasible solutions, the distributed solvers drive the primal residual down to the threshold and achieve an optimality gap below one percent, indicating that the distributed solutions are both feasible and near-optimal.
Case pglib_opf_case200_tamu
The system pglib_opf_case200_tamu is a 200-bus synthetic model representing a 230/115 kV network in Illinois. It was generated as part of the ARPA-E GRID DATA project and is designed to capture realistic transmission characteristics and operational patterns. The system is partitioned into three regions connected through ten tie-lines. The resulting multi-region topology is shown in Fig. 3.
Figure 3: Three-region topology for case200_tamu.
Figures 4 and 5 report the DC and AC primal residuals and optimality gaps for pglib_opf_case200_tamu.
Figure 4: primal residual (left) and optimality gap (right)
Figure 5: primal residual (left) and optimality gap (right)
Case pglib_opf_case300_ieee
The system pglib_opf_case300_ieee corresponds to the well-known IEEE 300-bus test case. Originally formulated in the IEEE Common Data Format and subsequently converted to MATPOWER, it has long served as a standard benchmark for power system optimization studies. The system is divided into three regions, with interconnections distributed across twelve tie-lines. The extracted topology is shown in Fig. 6.
Figure 6: Three-region topology for case300_ieee.
Figures 7 and 8 show the DC and AC verification results for pglib_opf_case300_ieee.
Figure 7: primal residual (left) and optimality gap (right)
Figure 8: primal residual (left) and optimality gap (right)
Case pglib_opf_case500_tamu
The system pglib_opf_case500_tamu is a 500-bus synthetic model of the South Carolina transmission grid, also produced under the ARPA-E GRID DATA initiative. It contains detailed line, transformer, and generator data reflecting real-world system behavior. The network is partitioned into five regions connected by sixteen tie-lines. The corresponding topology appears in Fig. 9.
Figure 9: Five-region topology for case500_tamu.
For pglib_opf_case500_tamu, Figures 10 and 11 illustrate DC and AC primal residuals and optimality gaps, respectively.
Figure 10: primal residual (left) and optimality gap (right)
Figure 11: primal residual (left) and optimality gap (right)
Case pglib_opf_case1354_pegase
The dataset pglib_opf_case1354_pegase is part of the PEGASE project, which produced large-scale synthetic European transmission networks intended for stability and optimization studies. This 1354-bus system includes high-density meshed areas characteristic of continental European grids. It is partitioned into eight regions with numerous inter-region connections reflecting the underlying network complexity. The multi-region layout is shown in Fig. 12.
Figure 12: Eight-region topology for case1354_pegase.
Figures 13 and 14 present the DC and AC verification curves for pglib_opf_case1354_pegase.
Figure 13: primal residual (left) and optimality gap (right)
Figure 14: primal residual (left) and optimality gap (right)
Case pglib_opf_case2869_pegase
The system pglib_opf_case2869_pegase represents a large portion of the European transmission network, consisting of 2869 buses, 510 generators, and 4582 branches. The partitioning procedure divides the system into twelve regions, interconnected through ninety-five tie-lines. The resulting multi-region topology is shown in Fig. 15. All buses, generators, and branches are fully accounted for across the regions, and the reference bus (Bus 1314) is assigned to Region 6.
Figure 15: Twelve-region topology for case2869_pegase.
For pglib_opf_case2869_pegase, Figures 16 and 17 illustrate that both DC and AC primal residuals converge to the threshold, while the optimality gaps remain below one percent.
Figure 16: Primal residual (left) and optimality gap (right)
Figure 17: Primal residual (left) and optimality gap (right)
Case pglib_opf_case4661_sdet
The system pglib_opf_case4661_sdet is a large-scale synthetic model developed using the Sustainable Data Evolution Technology (SDET) framework. With 4661 buses and dense connectivity, it reflects realistic large-area transmission dynamics. The system is separated into fourteen regions connected by 284 tie-lines. The partitioned structure is shown in Fig. 18.
Figure 18: Fourteen-region topology for case4661_sdet.
For pglib_opf_case4661_sdet, Figures 19 and 20 indicate that the DC and AC primal residuals converge to the threshold, with optimality gaps remaining below one percent.
Figure 19: primal residual (left) and optimality gap (right)
Figure 20: primal residual (left) and optimality gap (right)
Case pglib_opf_case9241_pegase
The largest benchmark system considered in DPLib is pglib_opf_case9241_pegase, a synthetic European transmission network derived from the PEGASE project. With 9241 buses and more than sixteen thousand branches, it reflects the characteristics of a highly interconnected continental-scale system. The partitioning procedure yields eighteen regions linked by 197 tie-lines. The resulting multi-region topology is shown in Fig. 21.
Figure 21: Eighteen-region topology for case9241_pegase.
For pglib_opf_case9241_pegase, Figures 22 and 23 confirm that the DC and AC primal residuals converge to the threshold, with optimality gaps remaining within one percent, even for this largest benchmark.
Figure 22: primal residual (left) and optimality gap (right)
Figure 23: primal residual (left) and optimality gap (right)
V Conclusion
This work presented DPLib, a reproducible and open-source benchmark library tailored for distributed and decentralized power system studies. DPLib addresses the gap in standardized datasets for distributed OPF research by offering a graph-based partitioning toolkit and over 20 multi-region test cases derived from MATPOWER. The library is complemented by modular ADMM-based DC and AC OPF solvers, which enable scalable validation across diverse case sizes. Numerical results demonstrate the accuracy of the generated test cases and convergence of the solvers, making DPLib a practical foundation for future research in scalable, privacy-preserving, and resilient power system analysis.
To support long-term usability, we are committed to maintaining DPLib as an actively developed open-source project. Future updates will include bug fixes, feature enhancements, expanded test systems, and improvements to the distributed OPF solvers as new algorithms become available. The DPLib repository will continue to be updated and versioned through GitHub, where users may report issues, request features, and contribute to the codebase.